### v1

@@ VII-B Discussion
To further explain the sensor overlapping results under the ${\tt oSCC}$ and ${\tt SCC}$ planners shown in Fig. 15, we look into the underlying differences in the planning algorithms. Overlapping occurs when adding minimum-cost connecting edges to the graphs to create an Eulerian graph, specifically when adding the minimum-cost path between the exit of one cell and the entry of the subsequent cell. It can also be caused by traversing an edge in the crack graph $\mathbb{G}_{\text{c}}$ more than once to achieve an optimum Eulerian path. All the connecting edges account for the overlapped sensing area and the increase in path length. The ${\tt oSCC}$ planner results in less sensor overlap and shorter path length than the ${\tt SCC}$ because ${\tt oSCC}$ reduces repeated connecting edges in the online planning process, especially when the cracks are congregate. In addition, comparing Figs. 15 with 15, the sensor overlapping and the path length of the uniformly distributed cracks are larger than those of the Gaussian distribution for both the ${\tt oSCC}$ and ${\tt SCC}$ planners. These observations confirm that the connecting edges are the main contributor to sensor overlapping because the uniform distribution of cracks results in their widespread scattering, which subsequently leads to an increase in the length of the connecting edges during the LP matching process in the algorithm.
We further conducted a study to understand how the crack dispersity impacts the ${\tt oSCC}$ planner’s performance. We used Gaussian distribution crack maps to illustrate the results. The crack dispersity is defined by the normal distribution of the crack centroid locations. As explained previously, three standard variations of the total workspace areas were used, namely, $\sigma_{1}=\frac{1}{12}$, $\sigma_{2}=\frac{1}{6}$, and $\sigma_{3}=\frac{1}{3}$. Fig. 16 shows the comparison of the computation time and robot path length under the ${\tt oSCC}$ and ${\tt GCC}$ planners with varying crack density and three dispersities. We chose the ${\tt GCC}$ planner as the benchmark as discussed above. Fig. 16 shows the computation time (total and per iteration) of the ${\tt oSCC}$ planner. The results show a similar trend for different crack dispersities. Fig. 16 presents a comparison of robot path length under three different levels of crack dispersity for both the ${\tt oSCC}$ and ${\tt GCC}$ planners. It is observed that the ${\tt oSCC}$ planner results in a reduction of robot path length when cracks are cluttered (i.e., for small values of $\sigma$). In contrast, when the crack maps are known a priori and the entire free space is not covered, the ${\tt GCC}$ planner does not exhibit a significant difference in robot path length. These observations align with the previously discussed relationship between crack dispersity and connecting edge length (i.e., the more scattered the cracks are, the longer the connecting edges), as well as the role of connecting edges between substantial cells in contributing to overlapping. The ${\tt GCC}$ planner only adds connecting edges between cracks and does not result in noticeable differences in path length for varying levels of crack dispersity. Therefore, connecting edges between substantial cells play a major role in increasing path length and overlapping.
While the robot’s ability to precisely optimize the path between adjacent cells is limited due to unknown cell dimensions, extensive experimentation and comparison with ${\tt GCC}$ and ${\tt SCC}$ planners show that the final path of the ${\tt oSCC}$ planner achieves a high level of efficiency in terms of travel distance. Statistical simulations and extensive experiment results in Table I and Figs. 14-16 confirm this efficiency.
@@
Figure 17: Robot path length with various crack detection sensor radii ($S$) under the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners for (a) uniformly distributed cracks and (b) Gaussian-based distributed cracks. The workspace $w=6.1$ m and the footprint radius $a=8.9$ cm are fixed, and the lines and shaded areas represent the mean and one standard deviation value, respectively.
Several limitations of this work can be further explored and improved in the future. The planning algorithms assumed that the nozzle motion was fast enough such that all targets within the robot footprint $\mathcal{F}$ could be covered in time. As a result, minimizing the robot’s traveling distance was considered as the objective, and this simplified treatment decoupled the planning of robot motion from nozzle movement. The experiments were conducted in a simulated indoor environment with drop cloths and paint to create cracks and filling action, as the primary focus of the work was on the motion planning algorithm development. The performance metrics were estimated using the paint width, and even though the selected cloth did not absorb paint heavily, the results were not perfectly accurate compared to real crack-filling. We only used one mobile robot to conduct sensing and footprint coverage in this study, and it would be interesting to extend the SIFC problem with multiple collaborative robots to increase efficiency. In this work, we made assumptions about known robot locations and constant crack widths to emphasize the core planning concept and the motion planning and control algorithmic development. For future work, our focus will be on addressing uncertainties in localization by employing advanced techniques in simultaneous localization and mapping and achieving real-time estimation of the crack characteristics. Additionally, we plan to integrate crack width considerations into the planning algorithm’s cost and explore adaptive control algorithms to adjust the robot’s footprint coverage strategy based on encountered crack characteristics. These enhancements will significantly improve the applicability of our approach, allowing it to handle further realistic scenarios.
@@
In this paper, we assume a known free space where the robot can move freely in any direction without obstacles. We utilize classic cell decomposition methods, specifically MCD, for both coverage tasks. The work in [19, 21] demonstrates the applicability of MCD to workspaces with various shapes and a finite number of obstacles, covering both known and unknown environments. Although obstacle handling is not the main focus of this paper, we acknowledge that in real-world situations, robot sensors can detect obstacles, and these obstacles can be geometrically represented in the environment using polygonal approximations. MCD can then incorporate the detected obstacles into the determination of cell boundaries to ensure that cells do not overlap with obstacles or cross their boundaries. The Reeb graph of the free space is constructed after cell decomposition, and it is combined with the crack graph. The robot’s motion planning algorithm, such as the ${\tt oSCC}$ planner, can generate a path to navigate through the cracks and cells while avoiding obstacles, ensuring complete coverage of the workspace. The work in [19] extended MCD into three dimensions (3D), enabling coverage of closed, orientable, and connected surfaces in 3D. While we showcase results in a known rectangular free space, the method’s adaptability permits its extension to free spaces with various shapes and even to uneven or vertical 3D surfaces, provided that the number of obstacles remains finite. This characteristic renders the method suitable for a wide range of real-world scenarios.
@@ VIII Conclusion
We have presented a motion planning and control design for simultaneous robotic sensor-based inspection and footprint coverage, with applications to crack detection and repair in civil infrastructure. To address the challenging task of simultaneously performing two complete coverage tasks in SIFC, we first proposed a graph-based target coverage algorithm for the mobile robot. Subsequently, we introduced a novel algorithm to solve the SIFC problem with unknown target information. This algorithm ensures the complete sensor scan of the workspace and the full footprint coverage of all targets, with near-optimal performance in terms of traveling distance. With the planned robot trajectory, the nozzle motion was coordinated to efficiently fill all cracks underneath the robot footprint. Extensive experimental results confirmed the effectiveness of the proposed motion planning and control algorithms under various target distributions. Furthermore, we discussed and demonstrated comparisons with other benchmark planning algorithms. The presented near-optimal and complete coverage planning algorithm has the potential to be applied to other robotic SIFC applications.
@@ Appendix A Sketch proof of Lemma 1
According to the assumption of the MCD, no two critical points change the slice connectivity at the same time. Thus, critical points that are collinear with the slice direction need special consideration. Let the endpoint node be a node with only one connected edge. With horizontal sweep direction, we treat vertical crack edges connecting with endpoint nodes as cell boundaries and those endpoint nodes as critical points, denoted as vertical critical points. For example, as shown in Fig. 18, $N_{1}$, $N_{2}$, and $N_{3}$ are vertical critical points. Since the vertical edge of the crack graph separates one cell into two adjacent cells, vertical critical points have two edges in the Reeb graph. One edge comes from the crack graph, and the other from one of the two adjacent cells. With this extension, the slice connectivity remains constant within each cell.
@@
Except for the vertical critical points, the remaining critical points of the target region are generated by the surface normal that is perpendicular to the sweep direction. Note that the target region is the “dilated” crack graph by a circular disk with a radius of $S$. If the boundary of the target region is convex, we always find a node in the crack graph within an $S$-distance of the critical point. As shown in Fig. 5, critical points $C_{2}$, $C_{4}$, and $C_{5}$ correspond to nodes $N_{1}$, $N_{5}$, and $N_{4}$ in the crack graph, respectively. If the boundary of the target region is concave, then the corresponding node in the crack graph is the intersection point of edges (intersection node), e.g., critical point $C_{3}$ is associated with node $N_{3}$ in Fig. 5. Therefore, every critical point on the boundary of the target region is associated with one node in the crack graph. This proves the lemma.
@@ Appendix B Sketch proof of Proposition 1
Because of the property of the MCD, all the critical points that are generated by the convex boundaries (i.e., convex critical points) are connected to three cells. Similarly, all the critical points that are generated by the concave boundaries (i.e., concave critical points) are connected to one cell. In the Reeb graph, the convex and concave critical points correspond to nodes of degree three and one, respectively. Let us denote the convex and concave critical points of the target region as $C_{\text{vex}}$ and $C_{\text{cav}}$ and their corresponding crack nodes by Lemma 1 as $N_{\text{vex}}$ and $N_{\text{cav}}$, respectively.
If $N_{\text{vex}}$ is an endpoint node, then it has degree one. If $N_{\text{vex}}$ is an intersection node, then its degree plus the number of concave critical points associated with $N_{\text{vex}}$ is odd. To form the Euler tour, the nodes with odd degrees need to be paired up with the least cost. Notice that the distance between $C_{\text{vex}}$ and $N_{\text{vex}}$ is the sensing range $S$. Because the distance between other nodes in the crack graph and $C_{\text{vex}}$ is greater or equal to $S$, connecting each pair of $C_{\text{vex}}$ to $N_{\text{vex}}$ results in parts of the minimum-cost Euler tour. For the case where the distance between two critical points $C_{1}$ and $C_{2}$ is less than $S$, as shown in Fig. 18, the distance between the corresponding nodes $N_{1}$ and $N_{2}$ in the crack graph is greater than $2S$. Therefore, it is optimal to combine the $C_{\text{vex}}$ (node of degree three) to their corresponding $N_{\text{vex}}$ (odd degree node) in the Euler tour. Those combined nodes have an even degree, guaranteeing that the robot is not stuck at such nodes.
For $C_{\text{cav}}$, $N_{\text{cav}}$ must be an intersection node. The parity of the intersection node of the crack graph is the same as the parity of the number of critical points associated with it. Therefore, pairing up these odd nodes with the least cost results in the optimal Euler tour. The vertical critical points are defined on the graph nodes by Lemma 1 and have a degree of two. The vertical critical points do not affect the optimality of the Euler tour. Thus, to find the minimum cost of the Euler tour, the edge of the crack graph never gets doubled, as all the critical points of the target region result in even degrees by connecting them to their corresponding nodes in the crack graph. To pair up other odd nodes in the Reeb graph, only edges corresponding to the cells (free space) get doubled. Doubling the selected edges means splitting the corresponding cells into two portions, which does not increase the cost of covering the whole area. Thus, as all parts of the free space and crack graph are covered exactly once, optimality is preserved. This proves the proposition.
@@ Appendix C Sketch proof of Proposition 2
The ${\tt oSCC}$ planner reinforces the connections of the critical points to their corresponding nodes in $\mathbb{G}_{\text{c}}$ until reaching its end. The resulting least-cost Euler tour, obtained from the doubling of selected edges, establishes an order in which the cells of the Reeb graph should be visited. The Euler tour ensures that no area is covered twice by not traversing any edge twice. Note that the covered spaces are removed from $\mathcal{W}$, and the cell coverage does not duplicate any covered area. Doubling the edges of cells does not increase any cost because the two cell components do not overlap. By definition, no edge of the Euler tour is traversed twice, and this implies no area is covered twice. Therefore, the algorithm ensures no redundancy in terms of individual cell coverage, as all free space is covered exactly once from the least-cost Euler tour, resulting in the most efficient path that minimizes the robot’s travel distance to locally cover all the individual cells when the connections of each covered cell are not considered.

### v2

@@ VII-B Discussion
To further explain the sensor overlapping results under the ${\tt oSCC}$ and ${\tt SCC}$ planners shown in Fig. 15, we look into the underlying differences in the planning algorithms. Overlapping occurs when adding minimum-cost connecting edges to create an Eulerian graph, specifically when adding the minimum-cost path between the exit of one cell and the entry of the subsequent cell. It can also be caused by traversing an edge in the crack graph $\mathbb{G}_{\text{c}}$ more than once to achieve an optimal Eulerian path. All connecting edges contribute to the overlapped sensing area and the increase in path length. The ${\tt oSCC}$ planner results in less sensor overlap and shorter path length than the ${\tt SCC}$ because ${\tt oSCC}$ reduces repeated connecting edges in the online planning process, especially when cracks are clustered. In addition, comparing Figs. 15 with 15, the sensor overlapping and the path length of the uniformly distributed cracks are larger than those of the Gaussian distribution for both the ${\tt oSCC}$ and ${\tt SCC}$ planners. These observations confirm that the connecting edges are the main contributor to sensor overlapping because the uniform distribution of cracks results in their widespread scattering, which subsequently leads to an increase in the length of the connecting edges during the LP matching process in the algorithm.
@@
We further conducted a study to understand how crack dispersity impacts the ${\tt oSCC}$ planner’s performance. We used Gaussian distribution crack maps to illustrate the results. The crack dispersity is defined by the normal distribution of the crack centroid locations. As explained previously, three standard variations of the total workspace areas were used, namely, $\sigma_{1}=\frac{1}{12}$, $\sigma_{2}=\frac{1}{6}$, and $\sigma_{3}=\frac{1}{3}$. Fig. 16 shows the comparison of the computation time and robot path length under the ${\tt oSCC}$ and ${\tt GCC}$ planners with varying crack densities and three dispersities. We chose the ${\tt GCC}$ planner as the benchmark for the reason discussed above. Fig. 16 displays the computation time (total and per iteration) of the ${\tt oSCC}$ planner. The results exhibit a similar trend for different crack dispersities. Fig. 16 presents a comparison of robot path length under three different levels of crack dispersity for both the ${\tt oSCC}$ and ${\tt GCC}$ planners. It is observed that the ${\tt oSCC}$ planner results in a reduction of robot path length when cracks are cluttered (i.e., for small values of $\sigma$). In contrast, when the crack maps are known a priori and the entire free space is not covered, the ${\tt GCC}$ planner does not exhibit a significant difference in robot path length. These observations align with the previously discussed relationship between crack dispersity and connecting edge length (i.e., the more scattered the cracks are, the longer the connecting edges), as well as the role of connecting edges between substantial cells in contributing to overlapping. The ${\tt GCC}$ planner only adds connecting edges between cracks and does not result in noticeable differences in path length for varying levels of crack dispersity. Therefore, connecting edges between substantial cells play a major role in increasing path length and overlapping. While the robot’s ability to precisely optimize the path between adjacent cells is limited due to unknown cell dimensions, extensive experimentation and comparison with ${\tt GCC}$ and ${\tt SCC}$ planners show that the final path of the ${\tt oSCC}$ planner achieves a high level of efficiency in terms of travel distance. Statistical simulations and extensive experimental results in Table I and Figs. 14-16 confirm this efficiency.
@@
Figure 17: Robot path length with various crack detection sensor radii ($S$) under the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners for (a) uniformly distributed cracks and (b) Gaussian-based distributed cracks. The workspace width $w=6.1$ m and the footprint radius $a=8.9$ cm are fixed, and the lines and shaded areas represent the mean and one standard deviation, respectively.
@@
In this paper, we assume a known free space where the robot can move freely in any direction without obstacles. We used classic cell decomposition methods, specifically MCD, for both coverage tasks. The work in [19, 21] demonstrates the applicability of MCD to workspaces with various shapes and a finite number of obstacles, covering both known and unknown environments. Although obstacle handling is not the main focus of this paper, we acknowledge that in real-world situations, robot sensors can detect obstacles, and these obstacles can be geometrically represented in the environment using polygonal approximations. MCD can then incorporate the detected obstacles into the determination of cell boundaries to ensure that cells do not overlap with obstacles or cross their boundaries. The robot’s motion planning algorithm, such as the ${\tt oSCC}$ planner, can generate a path to navigate through the cracks and cells while avoiding obstacles, ensuring complete coverage of the workspace. The work in [19] extended MCD into three dimensions (3D), enabling coverage of closed, orientable, and connected surfaces in 3D. While we showcase results in a known rectangular free space, the method’s adaptability permits its extension to free spaces with various shapes and even to uneven or vertical 3D surfaces, provided the number of obstacles remains finite. This characteristic renders the method suitable for a wide range of real-world scenarios.
Several limitations of this work can be further explored and improved in the future. The planning algorithms assumed that the nozzle motion was fast enough such that all targets within the robot footprint could be covered in time. As a result, minimizing the robot’s traveling distance was considered as the objective, and this simplified treatment decoupled the planning of robot motion from nozzle movement. The experiments were conducted in a simulated indoor environment with drop cloths and paint to create cracks and filling action, as the primary focus of the work was on motion planning algorithm development. The performance metrics were estimated using the paint width, and even though the selected cloth did not absorb paint heavily, the results were not perfectly accurate compared to real crack-filling. We only used one mobile robot to conduct sensing and footprint coverage in this study, and it would be interesting to extend the SIFC problem with multiple collaborative robots to increase efficiency. In this work, we made assumptions about known robot locations and constant crack width to emphasize the core planning concept and motion planning and control algorithmic development. For future work, our focus will be on addressing uncertainties in localization by employing advanced techniques in simultaneous localization and mapping and achieving real-time estimation of the crack characteristics. Additionally, we plan to integrate crack width considerations into the planning algorithm’s cost and explore adaptive control algorithms to adjust the robot’s footprint coverage strategy based on encountered crack characteristics. These enhancements will significantly improve the applicability of our approach, allowing it to handle further realistic scenarios.
@@ VIII Conclusion
We have presented a motion planning and control design for simultaneous robotic sensor-based inspection and footprint coverage, with applications to crack detection and repair in civil infrastructure. To address the challenging task of simultaneously performing two complete coverage tasks in SIFC, we first proposed a graph-based target coverage algorithm for the mobile robot. Subsequently, we introduced a novel algorithm to solve the SIFC problem with unknown target information. This algorithm ensured the complete sensor scan of the workspace and the full footprint coverage of all targets, with near-optimal performance in terms of traveling distance. With the planned robot trajectory, the nozzle motion was coordinated to efficiently fill all cracks underneath the robot footprint. Extensive experimental results confirmed the effectiveness of the proposed motion planning and control algorithms under various target distributions. Furthermore, we discussed and demonstrated comparisons with other benchmark planning algorithms. The presented near-optimal and complete coverage planning algorithm has the potential to be used to other robotic SIFC applications.
@@ Appendix A Sketch proof of Lemma 1
According to the assumption of MCD, no two critical points change the slice connectivity at the same time. Thus, critical points collinear with the slice direction require special consideration. Let the endpoint node be a node with only one connected edge. When using a horizontal sweep direction, vertical crack edges connecting with endpoint nodes are treated as cell boundaries, and those endpoint nodes are considered critical points, denoted as vertical critical points. For example, as illustrated in Fig. 18, $N_{1}$, $N_{2}$, and $N_{3}$ represent vertical critical points. Since the vertical edge of the crack graph divides one cell into two adjacent cells, vertical critical points have two edges in the Reeb graph. One edge originates from the crack graph, and the other from one of the two adjacent cells. This extension ensures that the slice connectivity remains constant within each cell.
@@
Except for the vertical critical points, the remaining critical points of the target region are generated by the surface normal perpendicular to the sweep direction. Note that the target region is the “dilated” crack graph by a circular disk with a radius of $S$. If the boundary of the target region is convex, we always find a node in the crack graph within an $S$-distance of the critical point. As shown in Fig. 5, critical points $C_{2}$, $C_{4}$, and $C_{5}$ correspond to nodes $N_{1}$, $N_{5}$, and $N_{4}$ in the crack graph, respectively. If the boundary of the target region is concave, then the corresponding node in the crack graph is the intersection point of edges (intersection node), e.g., critical point $C_{3}$ is associated with node $N_{3}$ in Fig. 5. Thus, every critical point on the boundary of the target region is associated with one node in the crack graph. This proves the lemma.
@@ Appendix B Sketch proof of Proposition 1
Because of the property of MCD, all the critical points that are generated by the convex boundaries (i.e., convex critical points) are connected to three cells. Similarly, all the critical points that are generated by the concave boundaries (i.e., concave critical points) are connected to one cell. In the Reeb graph, the convex and concave critical points correspond to nodes of degree three and one, respectively. Let us denote the convex and concave critical points of the target region as $C_{\text{vex}}$ and $C_{\text{cav}}$ and their corresponding crack nodes as $N_{\text{vex}}$ and $N_{\text{cav}}$, respectively, by Lemma 1.
If $N_{\text{vex}}$ is an endpoint node, then it has degree one. If $N_{\text{vex}}$ is an intersection node, then its degree plus the number of concave critical points associated with $N_{\text{vex}}$ is odd. To form the Euler tour, the nodes with odd degrees need to be paired up with the least cost. Notice that the distance between $C_{\text{vex}}$ and $N_{\text{vex}}$ is the sensing range $S$. Because the distance between other nodes in the crack graph and $C_{\text{vex}}$ is greater or equal to $S$, connecting each pair of $C_{\text{vex}}$ to $N_{\text{vex}}$ results in parts of the minimum-cost Euler tour. For the case where the distance between two critical points $C_{1}$ and $C_{2}$ is less than $S$, as shown in Fig. 18, the distance between the corresponding nodes $N_{1}$ and $N_{2}$ in the crack graph is greater than $2S$. Therefore, the optimal approach is to combine the $C_{\text{vex}}$ (node of degree three) with their corresponding $N_{\text{vex}}$ (odd degree node) in the Euler tour. These combined nodes have an even degree, guaranteeing that the robot is not stuck at such nodes.
For $C_{\text{cav}}$, $N_{\text{cav}}$ must be an intersection node. The parity of the intersection node of the crack graph is the same as the parity of the number of critical points associated with it. Therefore, pairing up these odd nodes with the least cost ensures the optimal Euler tour. The vertical critical points are defined on the graph nodes (by Lemma 1) and have a degree of two. The vertical critical points do not affect the optimality of the Euler tour. Thus, to find the minimum cost of the Euler tour, the edge of the crack graph never gets doubled, as all the critical points of the target region result in even degrees by connecting them to their corresponding nodes in the crack graph. To pair up other odd nodes in the Reeb graph, only edges corresponding to the cells (free space) get doubled. Doubling the selected edges means splitting the corresponding cells into two portions, which does not increase the cost of covering the whole area. Thus, as all parts of the free space and crack graph are covered exactly once, optimality is preserved. This proves the proposition.
@@ Appendix C Sketch proof of Proposition 2
The ${\tt oSCC}$ planner reinforces the connections of the critical points to their corresponding nodes in $\mathbb{G}_{\text{c}}$ until reaching its end. The resulting least-cost Euler tour, obtained from the doubling of selected edges, establishes an order in which the cells of the Reeb graph should be visited. The Euler tour ensures that no area is covered twice by not traversing any edge twice. Note that the covered spaces are removed from $\mathcal{W}$, and the cell coverage does not duplicate any covered area. Doubling the edges of cells does not increase any cost because the two cell components do not overlap. By definition, no edge of the Euler tour is traversed twice, and this implies no area is covered twice. Therefore, the algorithm ensures no redundancy in terms of individual cell coverage. All free space is covered exactly once from the least-cost Euler tour. This results in the most efficient path that minimizes the robot’s travel distance to locally cover all the individual cells when the connections of each covered cell are not considered.